# footnotetext: * Equal Contribution

### v1

@@ 4.4 Evaluation on the Kubric dataset
We report our quantitative 3D point tracking metrics in Table 6, and find that across all metrics, our method outperforms the baselines. Moreover, we find qualitatively that the optimized motion coefficients of the scene representation are coherently grouped with each moving object in the scene. We demonstrate this in Figure 6, where we show the first 3 PCA components of our optimized motion coefficients of evaluation scenes.
4.5 Ablation studies
We ablate various components of our method on the iPhone dataset in 3D tracking in Table  4.5. We first validate our choices of motion representation, namely the $\mathbb{SE}(3)$ motion bases parameterization, with two ablations: 1) “Per-Gaussian Transl.”: replacing our motion representation with naive per-Gaussian translational motion trajectories, and 2) “Transl Bases.”: keeping the motion bases representation but only using translational bases instead of $\mathbb{SE}(3)$. We find $\mathbb{SE}(3)$ bases lead to a noticable improvement in 3D tracking compared to both translational bases and per-Gaussian translational motion representation.
Table 3: Ablation Studies on iPhone dataset.
[…]
Ours (Full) 0.082 43.0 73.3
Transl. Bases 0.093 42.3 69.9
Per-Gaussian Transl. 0.087 41.2 69.2
No $\mathbb{SE}(3)$ Init. 0.111 39.3 65.7
No 2D Tracks 0.141 30.4 57.8
5 Discussion and Conclusion
Limitations. Although our approach demonstrates promising steps towards long-range 3D tracking and accurate reconstruction of complex 3D dynamic scenes, several limitations remain. Similar to most prior monocular dynamic view synthesis methods, our system still requires per-scene test-time optimization, which hinders its use in streamable applications. In addition, it cannot handle large changes in camera viewpoint, which require some kind of generative capabilities or data-driven priors to hallucinate unseen regions. Moreover, our approach relies on accurate camera parameters obtained from off-the-shelf algorithms, suggesting potential failure in scenes dominated by texture-less regions or moving objects. Finally, our method relies on user input to generate the mask of moving objects. A promising research direction would be to design a feed-forward network approach for jointly estimating camera poses, scene geometry, and motion trajectories from an unconstrained monocular video.
Please note that since we wrote this paper, there have been several concurrent works posted on arXiv [46, 95, 82, 54] that also address this setup of monocular 4D reconstruction from causal videos. As far as we know, all of them are optimization-based approaches, also utilizing strong off-the-shelf data-driven priors. We leave it to future work to compare these approaches.
Conclusion. We presented a new method for joint long-range 3D tracking and novel view synthesis from dynamic scene captured by a monocular video. Our approach models dense dynamic scene elements with a global set of 3D Gaussians that translate and rotate over time. We regularize the full-length 3D motion trajectories of each Gaussian, ensuring smoothness and low dimensionality, by parameterizing them as linear combinations of a compact set of rigid motion bases. To overcome the ill-posed nature of this problem in monocular capture settings, we further design the model to regularize by consolidating noisy data-driven observations into a globally consistent estimate of scene appearance, geometry, and motion. Extensive evaluations on both synthetic and real benchmarks demonstrate that our approach significantly improves upon prior state-of-the-art methods in both 2D/3D long-range tracking and novel view synthesis tasks.
Acknowledgement. We thank Ruilong Li, Noah Snavely, Brent Yi and Aleksander Holynski for helpful discussion. This project is supported in part by DARPA No. HR001123C0021. and IARPA DOI/IBC No. 140D0423C0035. The views and conclusions contained herein are those of the authors and do not represent the official policies or endorsements of these institutions.

### v2

@@ 4.4 Evaluation on Kubric Dataset
Table 2: 3D Tracking evaluation on Kubric dataset.
We report our quantitative 3D point tracking metrics in Table 2, and find that across all metrics, our method outperforms the baselines. Please see the supplement for figures of the optimized motion coefficients, which coherently group moving objects.
4.5 Evaluation on NVIDIA Dataset
We conduct experiments on seven scenes from the NVIDIA dataset [125], following the evaluation protocol of Dynamic Gaussian Marbles [94], which evaluates the task from a single static camera instead of interleaving different camera as discussed in  [25]. We adopt the same image resolution, camera parameters and depths maps as input. Please see the supplement for more details. For each scene, we compute covisibility masks between each evaluation view and the training view at each time step to exclude unobservable regions during evaluation. Results are reported in Tab. 3.
Method Balloon1 Balloon2 Jumping Truck
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
DGM [94] 23.83 0.81 0.066 23.62 0.81 0.088 20.02 0.73 0.134 27.31 0.86 0.055
Ours 23.90 0.79 0.058 23.91 0.83 0.074 20.08 0.69 0.151 27.65 0.88 0.048
Method Playground Umbrella Skating Mean
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
DGM [94] 16.68 0.60 0.177 24.99 0.70 0.083 27.79 0.90 0.052 23.46 0.77 0.093
Ours 16.87 0.55 0.168 23.27 0.57 0.167 27.91 0.92 0.047 23.37 0.75 0.102
Table 3: Evaluation on NVIDIA dataset. Our method is comparable with Dynamic Gaussian Marbles (DGM) [94].
4.6 Ablation studies
We ablate various components of our method on the iPhone dataset in 3D tracking in Tab. 4. We first validate our choices of motion representation, namely the $\mathbb{SE}(3)$ motion bases parameterization, with three ablations: 1) “Per-Gaussian Transl.”: replacing our motion representation with naive per-Gaussian translational motion trajectories, 2) “Per-Gaussian $\mathbb{SE}(3)$”: replacing our motion representation with naive per-Gaussian $\mathbb{SE}(3)$ motion trajectories, and 2) “Transl. Bases”: keeping the motion bases representation but only using translational bases instead of $\mathbb{SE}(3)$. We find that $\mathbb{SE}(3)$ bases significantly improve 3D tracking over both translational bases and per-Gaussian translational motion. It also outperforms per-Gaussian $\mathbb{SE}(3)$ in overall accuracy and offers noticeably better visual quality, as the latter exhibits significantly more popping and jitter artifacts.
[…]
Ours (Full) 0.082 43.0 73.3
Transl. Bases 0.093 42.3 69.9
Per-Gaussian $\mathbb{SE}(3)$ 0.083 43.6 70.2
Per-Gaussian Transl. 0.087 41.2 69.2
No $\mathbb{SE}(3)$ Init. 0.111 39.3 65.7
No 2D Tracks 0.141 30.4 57.8
Table 4: Ablation Studies on iPhone dataset.
5 Discussion
Limitations. Like most prior monocular view synthesis methods, it still requires per-scene test-time optimization, hindering streamable applications. Recent feed-forward methods for joint reconstruction and tracking  [22, 114, 116] offer a promising alternative, showing encouraging progress toward overcoming this limitation. Our method relies on off-the-shelf predictions for camera poses, geometry, and motion, which may degrade in textureless regions or under large motions. However, it also naturally benefits from advances in these components. Finally, it also requires user input to mask moving objects. Recent advances in moving object segmentation [27, 35] offer promising directions beyond our current manual solution and could be integrated into our pipeline.
Conclusion. We introduce a method for joint long-range 3D tracking and novel view synthesis from monocular video, representing dynamic elements via global 3D Gaussians with time-varying translations and rotations. Motion trajectories are regularized using low-dimensional rigid motion bases. By consolidating noisy observations into globally consistent scene estimates, our approach outperforms state-of-the-art methods in 2D/3D tracking and novel view synthesis across synthetic and real benchmarks.
Acknowledgement. We thank Ruilong Li, Noah Snavely, Brent Yi and Aleksander Holynski for helpful discussion. We are in memory of our beloved cat Sriracha, who will always be missed and loved. This project is supported in part by DARPA No. HR001123C0021. and IARPA DOI/IBC No. 140D0423C0035. The views and conclusions contained herein are those of the authors and do not represent the official policies or endorsements of these institutions.
Appendix
Appendix A Additional Preprocessing Details
Obtaining Camera Poses. Our method takes video sequences with known camera poses as input. To obtain camera poses for in-the-wild videos with moving objects, we adopt one of these two approaches depending on the type of input camera motion: (1) if there is sufficient camera motion parallax, we use COLMAP [88]’s SfM pipeline to obtain the camera poses and the sparse point clouds for the static regions, where we exclude keypoints in the foreground masks during the feature extraction stage. The foreground masks are generated using Track-Anything [120], a flexible and interactive tool for video object tracking and segmentation. The static point clouds produced by COLMAP [88] can then be used for aligning the affine-invariant monocular depth maps from Depth Anything [121]. (2) If the video is captured by a roughly stationary camera (small interframe camera baseline), COLMAP tends to fail catastrophically. Initial results on DAVIS use camera poses estimated with Unidepth [78] and DroidSLAM [98]. Specifically, we first employ Unidepth to predict metric depth and camera intrinsics. Using these predictions, we then estimate camera poses with DroidSLAM. (3) Camera parameters of in-the-wild videos in supplemental videos are estimated with MegaSaM [62].
Aligning Monocular Depth Maps. The dispairty output from Depth Anything [121] model is affine-invariant, so we need to align them with the cameras and reconstruction. To do so, we solve for a per-frame global scale and shift parameters that minimizes $\ell_{2}$ distance between the monocular disparity from Depth Anything and the disparity derived from SfM sparse point clouds or depth outputs from DroidSLAM or MegaSaM.
Computing Long-Range 2D Tracks. We utilize TAPIR [16] to compute long-range 2D tracks for a video. While the ideal scenario would involve computing full-length tracks for every pixel in every frame (i.e., exhaustive pairs of correspondences), this approach is prohibitively computationally expensive. We therefore only compute full-length tracks for pixels located on a grid for foreground moving objects in each frame. In our experiment, we set grid interval to $4$ (i.e., sampling a query point every 4 pixels). Due to our low-dimensional motion representation, we observe that our method can effectively operate with semi-dense 2D tracks without significant performance degradation During training, we filter out correspondences that TAPIR predicts with high uncertainty or those that are occluded.
Appendix B Additional Training Details
B.1 Initialization Details
During initialization stage, we first solve a Procrustes alignment problem for each cluster $b$, where we estimate $\mathbb{SE}(3)$ transformation between point sets $\{\mathbf{X}_{0}\}_{b}$ and $\{\mathbf{X}_{\tau}\}_{b}$ for all $\tau=0,\dots,T$ . We exclude point pair when either one of them is occluded, and weight the point pair using the uncertainty score predicted by TAPIR [16] when solving Procrustes. This process produces the initialization for the set of basis functions $\{\mathbf{T}_{0\rightarrow t}^{(b)}\}_{b=1}^{B}$. To initialize the motion coefficient $\mathbf{w}^{(b)}$ for each track, we compute the distance between the 3D location of the track in the canonical frame and the 3D location of each cluster center, and initialize each of the corresponding motion coefficient value to be exponentially decay with the distance.
We then optimize the $\boldsymbol{\mu}_{0}$, $\mathbf{w}^{(b)}$, and set of basis functions $\{\mathbf{T}_{0\rightarrow t}^{(b)}\}_{b=1}^{B}$ to fit the observed 3D tracks lifted by monocular depths and 2D tracks. Specifically, we enforce an $\ell_{1}$ loss between each of our predicted 3D track and its corresponding observed 3D track. In addition, we enforce the motion bases to be temporally smooth by adding an $\ell_{2}$ regularization on the acceleration of both the quaternion and the translation vector. We optimize the parameters using Adam [48] optimizer for 22k steps, where the initial learning rates for $\boldsymbol{\mu}_{0}$, $\mathbf{w}^{(b)}$, and $\{\mathbf{T}_{0\rightarrow t}^{(b)}\}_{b=1}^{B}$ are $\times 10^{-3}, 1×1$$\times 10^{-2}$ and $1\times 10^{-2}$, respectively. All learning rates are exponentially decayed to $\frac{1}{10}$ of their initial values during the optimization process.
B.2 Training Details
Gaussian Initialization. The aforementioned initialization stage gives the initialization of the mean of each Gaussian in the canonical frame. We follow the original 3D-GS [47] paper to initialize the scale $\mathbf{s}$, rotation $\mathbf{R}_{0}$, and opacity $o$ of each Gaussian in the canonical frame. The color $\mathbf{c}$ of each Gaussian is initialized as the pixel color at the projected location in the canonical frame.
Optimization. We use Adam [48] Optimizer to optimize all scene parameters. The learning rates for Gaussian’s canonical mean  $\mathbf{\mu}_{0}$, opacity $o$, scale $\mathbf{s}$, rotation $\mathbf{R}_{0}$ (parameterized as quaternion) and color $\mathbf{c}$ are set to $6\times 10^{-4}, 1×1$$\times 10^{-2}$, $5\times 10^{-3}$, $1\times 10^{-3}$, and $1\times 10^{-2}$, respectively. The learning rates for the $\mathbb{SE}(3)$ motion bases and the motion coefficients are set to $1.6\times 10^{-4}$ and $1\times 10^{-2}$ respectively. During each training iteration, we randomly select a batch of 8 query frames. For each query frame, we render the color, mask, depth, and the 3D track locations for 4 randomly selected target frames.
Loss weights and details. Our first set of loss function enforces our rendered results to match the per-frame pixelwise color, depth, and masks inputs. The coefficients for depth loss $\lambda_{\text{depth}}$ and mask loss $\lambda_{\text{mask}}$ are set to $0.5$ and $1.0$, respectively. We additionally add regularization to the estimated surface geometry via a depth gradient loss and per-Gaussian scale penalty. In particular, We add a pixelwise $\ell_{1}$ loss of spatial gradient between the depth renderings $\hat{\mathbf{D}}_{t}$ and corresponding reference depth maps $\mathbf{D}$, with a weight set to 1.0. In addition, we enforce the foreground Gaussian to be isotropic by incorporating a regularization term that penalizes standard deviations of the scale $\mathbf{s}$ along all three axes.
Our second set of losses supervise the motion of the Gaussian. The weights for the 2D tracking loss $\lambda_{\text{track-2d}}$ and the track depth loss $\lambda_{\text{track-depth}}$ are set to $2.0$ and $0.1$, respectively. The $L_{\text{track-2d}}$ is applied on normalized pixel coordinates (i.e., divided by the maximum image edge length). For the rigidity loss $L_{\text{rigidity}}$, we use foreground part masks from SAM automatic segmentation independently on each frame. For each training iteration, we sample 4 part masks for each query frame. For each mask, we sample 32 center points and find their 16 nearest neighbors within the mask. We compute the 3D track locations for all point samples in 4 target frames, and regularize the distances between the center points and their neighbors in the target frames to be similar to their distances in the query frame. We let $\lambda_{\textrm{rigidity}}=0.1$. We weight the neighbor distances with an exponential kernel $\exp(-\beta\|x-\bar{x}\|)$ with $\beta=2$, where $\bar{x}$ is the center point and $x$ is one of its neighbors. We additionally add motion smoothness regularization that enforces an $\ell_{2}$ loss on acceleration of the motion translation bases and motion quaternion bases, with a weight set to 0.1 and an $\ell_{2}$ loss on the acceleration along $z$-axis (in the camera frame) of the $\mu_{t}$, both through finite difference approximation.
Training with 2D Gaussian Splatting (2DGS) To enhance scene geometry estimation, we replace 3D Gaussian Splatting with 2D Gaussian Splatting [34]. Specifically, we employ an off-the-shelf monocular normal estimator [31] to generate monocular normal maps, $\tilde{N}$, which are then used to supervise both the rendered normals and the normals derived from the rendered depth. The supervision is achieved through the following losses:
[…]
$L_{N}=1-N^{T}\tilde{N}$ (12)
where the summation is over the splats intersect the current ray, $\omega_{i}$ represents the blending weight for $i$-th splat, and $n_{i}$ denotes the normal of $i$-th splat. Here, $N$ represents the normal derived from rendered depth map. This supervision ensures that the orientation and depth of the 2D splats are aligned with the monocular normal prediction.
Appendix C Additional Evaluation Details
In our evaluation, since the synthetic Kubric dataset [28] comes with groundtruth camera poses, we directly use the groundtruth camera poses for our experiments. For iPhone dataset [25], we observed that the provided camera poses from ARKit are not accurate, we thus perform an additional global bundle adjustment using COLMAP [88] to refine the camera poses while fixing the camera intrinsics. To maintain metric scale after refinement, we compute a global $\mathbb{SIM(}3)$ transformation for each scene to align the refined camera poses with the original metric-scale camera poses. This allows us to evaluate 3D tracking performance in metric scale.
Appendix D Visualization of Kubric Experiment
We find qualitatively that the optimized motion coefficients of the scene representation are coherently grouped with each moving object in the scene. We demonstrate this in Figure 6, where we show the first 3 PCA components of our optimized motion coefficients of evaluation scenes.
@@ Appendix D Visualization of Kubric Experiment
Appendix E Visualization of Complex Dynamic Scene
Our method is able to handle scenes with multiple moving objects. We visualize novel views and motion PCA of a real world complex scene in Fig. 7.
Figure 7: Novel view and motion coefficient PCA visualizations at time steps 0 and 54 of the school-girl sequence from the DAVIS dataset.
Appendix F NVIDIA Dataset Evaluation
We conduct experiments on seven scenes from the NVIDIA Dynamic Scenes dataset [125], following the evaluation protocol of Dynamic Gaussian Marbles [94]. Specifically, we use video footage from the static camera 4 for training and employ videos from cameras 3, 5, and 6 for evaluation. For consistency with Dynamic Gaussian Marbles [94]’s experiments, we use images at half-resolution for all seven experiments, though our method is compatible with high-resolution images as well. Camera poses are estimated using COLMAP [88], and depths are predicted with Depth Anything [121] and subsequently aligned with the COLMAP point cloud. The quantitative results reported in this paper differ from those in DGM [94]. Specifically, we compute covisibility masks between training and test views and apply them during evaluation, whereas DGM [94] inpaints unobserved regions in test views for evaluation. This difference in evaluation procedure accounts for the variation in reported performance. We will make the covisibility masks publicly available to facilitate future research.
Appendix G DynMF Implementation Detail
DynMF  [53] shares similar design choices with our method. Since they don’t have public code available, we implement our own version of it and provide implementation details here. Follow equation (4) in DynMF [53], each motion basis is represented as a MLP:
$b_{j}(t)=MLP_{j}\left(\frac{t}{T}\right)$ (13)
where $b_{j}$ is the jthj-th motion basis, $t$ is the queried time step, and $T$ is the number of time frames. We use a total of 10 motion bases. We apply positional encoding with 10 frequency bandwidths [71] to the input $\frac{t}{T}$. Following DynMF [53], we apply displacement motion to Gaussian’s means and rotation quaternions as follows:
$\mu(t)=\mu_{c}+\sum_{j=1}^{10}c_{j}b_{j}^{\mu}(t)$ (14)
where $\mu_{c}$ is the Gaussian means in the canonical frame and $\{b_{j}^{\mu}(t)\}_{j=1}^{10}$ are the mean motion bases represented as MLPs.
and
$q(t)=q_{c}+\sum_{j=1}^{10}c_{j}b_{j}^{R}(t)$ (15)
where $q_{c}$ is the Gaussian rotation represented as quaternion in the canonical frame and $\{b_{j}^{R}(t)\}_{j=1}^{10}$ are the rotation motion bases represented as MLPs.The motion coefficients $\{c_{j}\}_{j=1}^{10}$ are shared between mean and rotation motions.